A North Face retail store in Chicago sells 120 jackets per week. Each jacket costs the store $100 and the company has an annual holding cost of 25 percent. The fixed cost of a replenishment order (including transportation) is $100. The store currently places a replenishment order every month for 500 jackets. What is the annual holding cost? What is the annual ordering cost? What is the annual total cost including the purchasing cost? If the retail store wants to minimize the annual total inventory cost, what order size do you recommend? How much would the optimal order quantity reduce the annual total inventory cost relative to the current policy?
Answer #1

Annual Holding Cost = 25*500/2

Annual Ordering Cost = 100*12 = 1200

Annual Total Cost = Purchase Cost+Holding Cost+Ordering Cost

= 500*12*100 + 6250+1200

607470

EOQ ={ 2(100)(6000)/25} raised to power ½

= 219 .08 units or 219 units

Total Annual Inventory Cost = 600000+28*100+25*219/2

= 605,537.5

Savings in Cost = 1932.5